OpenStudy (anonymous):

what is the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -x + 5?

OpenStudy (angela210793):

Do u have any idea how to do this??

OpenStudy (anonymous):

supossed to say y=-1/2x+5 sorry :p

OpenStudy (anonymous):

I wouldn't know where to begin

OpenStudy (angela210793):

the slope of the line, perpendicular to the one you're given is \[slope=-1/m\] (m is the slope of the line ur given) do u know what is the slope of y=-x+5?

OpenStudy (anonymous):

yes we are given the equation y=-1/2x+5 and using point slope which I belive is y=mx+b then according to that 1/2 would be the slope

OpenStudy (angela210793):

ur almost right..the slope is -1/2(negative 1/2) so the slope of the line ur trying to find is -1/(-1/2)=2 the equation of the line when ur given the slope and a point is :\[y-y0=m(x-x0)\]

OpenStudy (angela210793):

x0=4 and y0=-6....can u do the substituition???

OpenStudy (anonymous):

D: what am I substituting for?

OpenStudy (angela210793):

Sorry...I've been doing too much maths lately :S